# Existence, uniqueness, and regularity of

### v1

@@ 1 Introduction and Main Result
Theorem 1 establishes the regularity of solutions to fully nonlinear parabolic obstacle problems when the operator is only required to be measurable and the obstacle is may be irregular as noted above. Classic references on obstacle and free boundary problems include Bensoussan and Lions, (1978), Friedman, (1982) and Petrosyan et al., (2012). When the obstacle is smooth enough and the operator $F$ is linear with smooth coefficients and smooth nonhomogeneous term (for example, if the obstacle is $\mathcal{C}^{1,2}$ and the coefficients and nonhomogeneous term are $\alpha$-Hölder continuous), the obstacle problem has a unique solution in $W^{1,2,\infty}\left(\bar{\mathcal{Y}}\right)$, see, e.g., Friedman, (1982). This result extends to the nonlinear case when both the operator and obstacle are smooth and the operator satisfies strong convexity and growth conditions (Petrosyan and Shahgholian,, 2007; Audrito and Kukuljan,, 2023). In these cases, the regularity of solutions to various types of obstacle problems has been extensively studied under appropriate regularity assumptions on the boundary of the domain, the obstacle, the coefficients of the linear operator, and the nonhomogeneous term.
For the case of irregular nonlinear operators, but without obstacles, solvability and regularity results for elliptic and parabolic equations have been obtained under either a small mean or small $L^{\infty}$ oscillation condition in Caffarelli, (1989), Escauriaza, (1993), Caffarelli et al., (1996), Winter, (2009), or Krylov, (2010) for the elliptic case, and in Crandall et al., (1998), Crandall et al., (1999), Crandall et al., (2000), Dong et al., (2013), and Krylov, (2017) for the parabolic case, among others. The recent monograph Krylov, (2018) offers an up-to-date general treatment.
The regularity of viscosity solutions for the parabolic obstacle problem is studied by (Byun et al.,, 2018) in the linear case with VMO leading coefficients when the obstacle is in $W^{1,2,p}$ and by (Byun et al.,, 2022) in the fully nonlinear elliptic case when the operator satisfies the same VMO assumption as in our study and the obstacle is in $W^{1,2,p}$.
Our approach builds on the approximation argument developed in Byun et al., (2018) and Byun et al., (2022). We extend the argument in Byun et al., (2018) to the fully nonlinear parabolic case and allow the operator to be only measurable rather than continuous and the obstacle to be irregular. We first obtain a $W^{1,2,p}$-estimate when the obstacle is in $W^{1,2,p}$ using known estimates on the regularity of solutions of the Dirichlet problem, e.g., Theorem 12.1.7 and 15.1.3 in Krylov, (2018). We then show by induction that the existence and regularity results for a single $W^{1,2,p}$-obstacle carry to the pointwise maximum of a finite number of $W^{1,2,p}$-obstacles (Lemma 2). Finally, we use a limit argument to consider the general case.
Our argument relies on PDE methods and generalizes existing results derived by probabilistic methods in the context of optimal stopping. [^5] The connection between obstacle problems and stochastic control and stopping problems and stochastic games had received considerable attention in the literature. See, e.g., (Bensoussan and Lions,, 1978; Friedman,, 1982; Karatzas and Sudderth,, 2001; Peskir and Shiryaev,, 2006; Strulovici and Szydlowski,, 2015). For example, Décamps et al., (2006) consider the elliptic obstacle equation associated with an optimal stopping problem when the payoff upon stopping (i.e., the obstacle) is the maximum of two smooth convex functions. The authors use a local time argument to show that the decision-maker never stops at the convex kink. This, in turn, implies that the value function, hence the solution of (1), is smooth. Theorem 1, together with a Sobolev embedding theorem, allows us to recover their result: It guarantees that the solution of (1) is continuously differentiable in space, and, thus, that the decision-maker never stops at a point of nondifferentiability of the stopping payoff.
Theorem 1 establishes the existence, uniqueness, and $W^{1,2,p}$-regularity of the solution to a fully nonlinear obstacle problem when the primitives are only measurable and the obstacle is the supremum of arbitrary functions in $W^{1,2,p}$. This generalization of earlier results is crucial to accommodate settings in which the operator or the payoff upon stopping may not be smooth. For example, relaxing the operator’s continuity is crucial to analyzing the Hamilton-Jacobi-Bellman equation associated with Markovian equilibria in stochastic games, where best response functions may be discontinuous (Kuvalekar and Lipnowski,, 2020). Such discontinuities are ubiquitous in games with finite action sets.
1.1 Outline of the Paper
Section 2 introduces some notation and definitions. Section 3 contains two results needed for the proof of Theorem 1: (i) A comparison principle that applies to fully nonlinear parabolic obstacle problems with measurable ingredients (Proposition 1), and (ii) A stability theorem for solutions to obstacle problems in $W^{1,2,p}\left(\mathcal{Y}\right)$ (Theorem 2). Section 4 establishes our main result when $A$ is finite and $\mathcal{X}$ is smooth. Section 5.1 proves Theorem 1, first considering smooth domains and general separable set $A$ , then by using an approximation argument to treat Lipschitz domains.
2 Preliminaries
2.1 Notation
  • For $D>0$, $B_{D}(x)$ denotes the open ball of radius $D$ centered around $x$, and $C_{D}(t,x)=[t,t+D)\times B_{D}(x)$.
  • $\mathcal{C}^{k}\left(\mathcal{Y}\right)$ is the space of $k$-times continuously differentiable functions on $\mathcal{Y}$.
  • $W^{1,2,p}\left(\mathcal{Y}\right)$ denotes the Sobolev space of functions whose first weak time derivative and second weak space derivative are $L^{p}$-integrable. $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ is the space of functions that belong to $W^{1,2,p}\left(\mathcal{Y}^{\prime}\right)$ for all compact subsets $\mathcal{Y}^{\prime}$ of $\mathcal{Y}$.
  • $u_{t},u_{x}$, and $u_{xx}$ stand for the first weak derivatives of $u$ with respect to $t$, the first weak derivatives of $u$ with respect to $x$, and the second weak derivatives of $u$ with respect to $x$, respectively.
[…]
  • For $k\in\mathbb{N}$, an open bounded subset of $\mathbb{R}^{d}$ is $\mathcal{C}^{k,Lip}$ if, for every $x\in\partial\mathcal{X}$, there exists a neighborhood $V$ of $x$, $\alpha>0,\beta>0$, an affine map $T:\mathbb{R}^{d-1}\times\mathbb{R}\to\mathbb{R}^{d}$, and a map $\phi:B_{\alpha}(0)\subseteq\mathbb{R}^{d-1}\to\mathbb{R}$ that is either Lipschitz continuous if $k=0$, or $k$-times continuously differentiable with Lipschitz derivatives for $k\geq 1$ such that
    $\partial\mathcal{X}\cap V=T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\eta=\phi(\xi)\right\}\right),$
    and
    $T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\phi(\xi)<\eta<\phi(\xi)+\beta\right\}\right)\subseteq\mathcal{X}\cap V$
    and $T\left(\left\{(\xi,\eta)\in\mathbb{R}^{d-1}\times\mathbb{R}\,:\,\xi\in B_{\alpha}(0)\text{ and }\phi(\xi)-\beta<\eta<\phi(\xi)\right\}\right)\subseteq\mathbb{R}^{d}\setminus\left(\bar{\mathcal{X}}\cap\bar{V}\right).$
    The first part definition requires that there locally exists a coordinate system such that the boundary coincides locally with the graph of a function whose $k^{th}$ derivative (or the function itself if $k=0$) is Lipschitz. The two inclusions in the second part of the definition guarantee that the interior and the exterior of the domain are nonempty and locally contain a cone around any point of the boundary. They rule out, for example, the domain $\{(x,y)\,:\,x>0,\,|y|<x^{2}\}$.
  • The $\mathcal{C}^{k,Lip}$-norm of the boundary of a $\mathcal{C}^{k,Lip}$ open bounded subset $\mathcal{X}$ is defined by $\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{k,Lip}}=\underset{x\in\partial\mathcal{X}}{\sup}\,\left\|\phi\right\|_{\mathcal{C}^{k,1}(V)}$ where $V$ and $\phi$ are as defined above.
  • We will often use the letter $C$ to denote bounds appearing in various estimates, and which can be explicitly computed in terms of primitives of the problem.
2.2 Solution concepts and definitions
Definition 1
A function $u$ is an $L^{p}$-subsolution (respectively, $L^{p}$-supersolution) of Problem (1) if the following conditions hold: (i) $u\in{W}^{1,2,p}_{loc}(\mathcal{Y})\cap\mathcal{C}^{0}(\bar{\mathcal{Y}})$, (ii) $u\leq b$ (respectively, $\geq b$) on $\partial{\mathcal{Y}}$, and (iii)
$\max\left\{\,u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\,\,(\text{respectively, }\leq 0)\text{ a.e. in }\mathcal{Y}_{T}.$
We will also use viscosity solutions in stating and applying the comparison principle of Section 3.1. There are several concepts of viscosity solutions depending on the choice of test functions used in the definition. We use test functions in $W^{1,2,p}\left(\mathcal{Y}\right)$, which corresponds to what is sometimes called “$L^{p}$-viscosity solutions.” [^6] When $F$ is continuous, one can take the test functions to be in $\mathcal{C}^{1,2}$. In this case, the solution is called a $\mathcal{C}$-viscosity solution, and $\mathcal{C}$- and viscosity solutions coincide (Lemma 2.9 in Crandall et al., (2000)).
Definition 2
A continuous function $u:\bar{\mathcal{Y}}\to\mathbb{R}$ is a viscosity subsolution (respectively, supersolution) of (1) if (i) $u\leq b$ (respectively, $\geq b$) on $\partial{\mathcal{Y}}_{T}$, and (ii) for all $(t_{0},x_{0})$ and all $\varphi\in{W}^{1,2,p}_{loc}(\mathcal{Y}_{T})$ such that $u-\varphi$ (respectively $\varphi-u$) has a maximum at $(t_{0},x_{0})$, one has
$\underset{(t,x)\to(t_{0},x_{0})}{\text{ess}\limsup}$ $\,\max\left\{\,\varphi_{t}+F\left(t,x,u,\varphi_{x},\varphi_{xx}\right),\,g-u\right\}\geq 0$
[…]
Lemma 1
Let $u$ be an $L^{p}$-subsolution (supersolution, respectively) of (1). Then $u$ is a viscosity subsolution (supersolution, respectively) of (1).
@@ 3.1 Comparison Principle
Comparison principles for $L^{p}$-solutions and viscosity solutions of the Dirichlet problem are well-known in the literature [^7] See, e.g., Proposition 2.10 in Crandall et al., (2000) under a slight strengthening of the structure condition (SC). [^8] In particular, previous results assume that $F$ is Lipschitz continuous in $u_{x}$ and nonincreasing in $u$. We extend these results to fit the assumptions of Theorem 3. Our comparison principle is valid for a large class of nonlinear parabolic obstacle problems as long as $F$ satisfies the structure condition (SC) and does not increase too fast with $u$ (as required by Assumption 8).
@@
  • Let $u$ be an $L^{p}$-subsolution and $v$ be an $L^{p}$-viscosity supersolution of
@@
    If $v\geq u$ on $\partial\mathcal{Y}$, then $v\geq u$ in $\bar{\mathcal{Y}}$.
  • Let $u$ be an $L^{p}$-supersolution and $v$ be an $L^{p}$-viscosity subsolution of
@@
    If $v\leq u$ on $\partial\mathcal{Y}$, then $v\leq u$ in $\bar{\mathcal{Y}}$.
@@
$\max\left\{\,u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0,$
such that $v\geq u$ on $\partial\mathcal{Y}$. We will prove that $v\geq u$ in $\bar{\mathcal{Y}}$.
Let $w=u-v$. Let =$\mathcal{E}=\left\{(t,x)\in\bar{\mathcal{Y}}\,:\,w>0\right\}$ and note that $\mathcal{E}\subset\mathcal{Y}$, since $w\leq 0$ on $\partial\mathcal{Y}$. So, the conclusion follows if $\mathcal{E}$ is empty.
Suppose for a contradiction that $\mathcal{E}\neq\emptyset$. Note that $w$ is continuous since both $u$ and $v$ are continuous on $\bar{\mathcal{Y}}$ by definition. Therefore, $\mathcal{E}$ is an open bounded domain.
We first show that $w$ is a viscosity subsolution of
$\begin{cases}\max\left\{\,w_{t}+H\left(t,x,w,w_{x},w_{xx}\right),\,-w\right\}=0\text{ in }\mathcal{Y}\\ w=0\text{ on }\partial\mathcal{Y},\end{cases}$ (5)
@@
To see this, we first note that $w=0$ on $\partial\mathcal{E}$, so condition (i) in the definition of viscosity subsolutions (here, $w\leq 0$) is satisfied. Next, let $\varphi\in W^{1,2,p}\left(\mathcal{Y}\right)$ be such that $w-\varphi$ has a maximum at $(t_{0},x_{0})$, i.e., $u-\varphi-v$ has a maximum at $(t_{0},x_{0})$. By construction, $u-\varphi\in W^{1,2,p}$ and $v$ is a viscosity supersolution of (3). Therefore,
$\underset{(t,x)\to(t_{0},x_{0})}{\text{ess}\liminf}$ $\,\max\left\{\,u_{t}-\varphi_{t}+F\left(t,x,u-w,u_{x}-\varphi_{x},u_{xx}-\varphi_{xx}\right),\,g-(u-w)\right\}\leq 0.$
It follows that
$\underset{(t,x)\to(t_{0},x_{0})}{\text{ess}\limsup}$ $\,\max\left\{\,\varphi_{t}+H\left(t,x,w,\varphi_{x},\varphi_{xx}\right),\,-w\right\}\geq 0.$
That is, $w$ is a viscosity subsolution of (5).
By the structure condition (SC), for any test function $\varphi\in W^{1,2,p}\left(\mathcal{Y}\right)$ and almost every $(t,x)\in\mathcal{Y}$,
@@
Therefore, $w$ is also a viscosity subsolution of
$\begin{cases}\max\left\{-w,\,w_{t}+\mathcal{P}_{\lambda,\Lambda}^{+}\left(w_{xx}\right)+\omega_{1}\left(\left|w\right|\right)+\omega_{2}\left(\right|w_{x}\left|\right)\right\}=0\text{ in }\mathcal{Y},\\ w=0\text{ on }\partial\mathcal{Y}.\end{cases}$
Therefore, $w$ is a viscosity subsolution of
$\begin{cases}w_{t}+\mathcal{P}_{\lambda,\Lambda}^{+}\left(w_{xx}\right)+\omega_{1}\left(\left|w\right|\right)+\omega_{2}\left(\right|w_{x}\left|\right)=0\text{ in }\mathcal{E},\\ w=0\text{ on }\partial\mathcal{E}.\end{cases}$
Note that 00 is a classical solution of the above equation, hence a (continuous) viscosity supersolution. Theorem 4.7 in Giga et al., (1991) then implies that $w\leq 0$ in $\mathcal{E}$: a contradiction. So $\mathcal{E}=\emptyset$, and $v\geq u$ in $\bar{\mathcal{Y}}$.
The proof when $u$ is an $L^{p}$-supersolution, $v$ is a viscosity subsolution, and $u\geq v$ on $\partial\mathcal{Y}$ follows the same steps. $\blacksquare$
@@
Proposition 1 also holds when $F$ is degenerate elliptic (i.e, $\lambda=0$ in Assumption 4) since Theorem 4.7 in Giga et al., (1991) applies to degenerate elliptic operators. However, the existence of $L^{p}$-sub- or $L^{p}$-supersolutions is not guaranteed in that case.
Proposition 1 combined with Lemma 1 yields the following corollary, which establishes the uniqueness of $L^{p}$-solutions.
@@
Suppose that Assumptions 4 and 8 hold. Let $u$ and $v$ be, respectively, an $L^{p}$-subsolution and an $L^{p}$-supersolution of
@@
If $v\geq u$ on $\partial\mathcal{Y}$, then $v\geq u$ in $\bar{\mathcal{Y}}$.
@@ 3.2 Stability Theorem
Stability theorems for both $L^{p}$-solutions and viscosity solutions of the Dirichlet problem without obstacle are well-known under a slight strengthening of the structure condition (SC) [^9] In particular, $F$ has to be Lipschitz continuous in $u_{x}$ and nonincreasing in $u$. , see, e.g., Theorem 6.1 in Crandall et al., (2000). We prove such a stability theorem for our setting, as it is essential to prove 1. This stability theorem is valid for a large class of nonlinear parabolic obstacle problems as long as $F$ satisfies the structure condition (SC).
@@
    for all $n\in\mathbb{N}$, $u^{n}=b^{n}$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}^{n}$, and $b^{n}\to b$ pointwise, with $b$ continuous on $\partial\mathcal{Y}$.
Let $F,F^{n}$, $n\in\mathbb{N}$, be operators defined on $\mathcal{Y}\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}^{d}$ that satisfy Assumption 4 uniformly in $n$. Let $g^{n}$, $n\in\mathbb{N}$, be continuous functions on $\bar{\mathcal{Y}}$, such that $g^{n}\leq b^{n}$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}^{n}$.
Suppose that, for any fixed $u_{xx}\in L^{p}\left(\mathcal{Y}\right)$, $F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. in $\mathcal{Y}$, and, either (i) $g^{n}\to g$ uniformly on compact subset of $\mathcal{Y}$, or (ii) $g^{n}\uparrow g$ pointwise on $\mathcal{Y}$. If $u^{n}$, $n\in\mathbb{N}$, is an $L^{p}$-subsolution (supersolution) of
$\begin{cases}\max\left\{u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right),\,g^{n}-u^{n}\right\}=0\text{ in }\mathcal{Y}^{n},\\ u^{n}=b^{n}\text{ on }\partial\mathcal{Y}^{n},\end{cases}$ (Eqn)
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ in }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (Eq)
@@
Note that, for all $n\in\mathbb{N}$, $u^{n}\geq b^{n}$ on $\partial\mathcal{Y}$. Taking limits, $u\geq b$ on $\partial\mathcal{Y}$.

### v2

@@ 3 Main Result
Corollary 1
Suppose that Assumptions 18 hold for some $p\in(d+2,\infty)$. Then, the unique $L^{p}$-solution $u$ of (1) is continuously differentiable with respect to the space variable $x$ on $\mathcal{Y}$.
For stopping problems, the solution of (1) coincides with the value function. Corollary 1 then (i) implies that smooth pasting in space holds, and (ii) confirms that the decision maker never stops at a kink of $g$, as shown by Décamps et al., (2006) in the specific problem of irreversible investment.
Proof.
By Theorem 1, the unique solution is in $W^{1,2,p}_{loc}(\mathcal{Y})$. By a Morrey-Sobolev embedding type theorem (Lemma 3.3, page 80 in Ladyženskaja et al., (1968)), $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ is embedded in $\mathcal{C}^{0,1}\left(\mathcal{Y}\right)$. ∎
@@ 4 Comparison and Stability for fully nonlinear parabolic obstacle problems with measurable ingredients
Section 4 contains two results needed for the proof of Theorem 1: (i) A comparison principle that applies to fully nonlinear parabolic obstacle problems with measurable ingredients (Proposition 1), and (ii) A stability theorem for solutions to obstacle problems in $W^{1,2,p}\left(\mathcal{Y}\right)$ (Theorem 2).
@@ 4.1 Comparison Principle
Comparison principles for $L^{p}$-solutions and viscosity solutions of the Dirichlet problem are well-known in the literature. We provide here a more general result that encompasses obstacle problems and includes operators that are less restrained than those considered in the literature. [^10] For instance, Proposition 2.10 in Crandall et al., (2000) gives a comparison principle when $F$ is Lipschitz continuous in $u_{x}$ and nonincreasing in $u$. Our comparison principle is valid for a large class of nonlinear parabolic obstacle problems as long as $F$ satisfies the structure condition (SC) and does not increase too fast with $u$ (as required by Assumption 8).
@@
  • Let $u$ be an $L^{p}$-subsolution [^11] Note that we do not impose here conditions on the functions at the boundary $\partial\mathcal{Y}$; that is, we dispense with requirement (ii) in Definitions 1 and 2 and allow $u$ and $v$ to take a priori any values on $\partial\mathcal{Y}$. This is because the comparison principle only depends on the condition that $v\geq u$ on $\partial\mathcal{Y}$, which can be used to satisfy Definitions 1 and 2, including requirement (ii), for functions that depend on $u$ and $v$ only through the difference $v-u$. and $v$ be a viscosity supersolution of
@@
    If $v\geq u$ on $\partial\mathcal{Y}$, then $v\geq u$ on $\bar{\mathcal{Y}}$.
  • Let $u$ be an $L^{p}$-supersolution [^12] The same observation regarding the absence of a constraint on $\partial\mathcal{Y}$ applies here as in the previous footnote. and $v$ be a viscosity subsolution of
@@
    If $v\leq u$ on $\partial\mathcal{Y}$, then $v\leq u$ on $\bar{\mathcal{Y}}$.
@@
$\max\left\{\,u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0$
such that $v\geq u$ on $\partial\mathcal{Y}$. We will prove that $v\geq u$ on $\bar{\mathcal{Y}}$.
Let $w=u-v$ and =$\mathcal{E}=\left\{(t,x)\in\bar{\mathcal{Y}}\,:\,w>0\right\}$. We need to show that $\mathcal{E}$ is empty. Suppose by way of contradiction that it is not. We note that $\mathcal{E}\subset\mathcal{Y}$ since $w\leq 0$ on $\partial\mathcal{Y}$. Moreover, $w$ is continuous since both $u$ and $v$ are by construction continuous on $\bar{\mathcal{Y}}$ Therefore, $\mathcal{E}$ is an open bounded domain.
We begin by showing that $w$ is a viscosity subsolution of the problem
$\begin{cases}\max\left\{\,w_{t}+H\left(t,x,w,w_{x},w_{xx}\right),\,-w\right\}=0\text{ on }\mathcal{E}\\ w=0\text{ on }\partial\mathcal{E}\end{cases}$ (5)
@@
To see this, we first note that $w=0$ on $\partial\mathcal{E}$, so condition (ii) in the definition of viscosity subsolutions (which, here, amounts to $w\leq 0$) is satisfied. Next, let $\varphi\in W^{1,2,p}\left(\mathcal{E}\right)$ be such that $w-\varphi$ has a maximum at $(t_{0},x_{0})$ in $\mathcal{E}$, i.e., $u-\varphi-v$ has a maximum at $(t_{0},x_{0})$. By construction, $u-\varphi\in W^{1,2,p}_{loc}\left(\mathcal{E}\right)$ and $v$ is a viscosity supersolution of (3). Therefore,
$\underset{(t,x)\to(t_{0},x_{0})}{\mathrm{ess}\liminf}$ $\,\max\left\{\,u_{t}-\varphi_{t}+F\left(t,x,u-w,u_{x}-\varphi_{x},u_{xx}-\varphi_{xx}\right),\,g-(u-w)\right\}\leq 0.$
Combining this result with the fact that $u$ is an $L^{p}$-subsolution of (3) then yields
$\underset{(t,x)\to(t_{0},x_{0})}{\mathrm{ess}\limsup}$ $\,\max\left\{\,\varphi_{t}+H\left(t,x,w,\varphi_{x},\varphi_{xx}\right),\,-w\right\}\geq 0.$
This shows that $w$ is a viscosity subsolution of (5).
By the structure condition (SC), for any test function $\varphi\in W^{1,2,p}\left(\mathcal{E}\right)$ and almost every $(t,x)\in\mathcal{E}$,
@@
Therefore, $w$ is also a viscosity subsolution of the problem
$\begin{cases}\max\left\{-w,\,w_{t}+\mathcal{P}_{\lambda,\Lambda}^{+}\left(w_{xx}\right)+\omega_{1}\left(\left|w\right|\right)+\omega_{2}\left(\right|w_{x}\left|\right)\right\}=0\text{ on }\mathcal{E},\\ w=0\text{ on }\partial\mathcal{E}.\end{cases}$
By construction of the domain $\mathcal{E}$ and our earlier observation that $w=0$ on $\partial\mathcal{E}$, this implies that $w$ is a viscosity subsolution of
$\begin{cases}w_{t}+\mathcal{P}_{\lambda,\Lambda}^{+}\left(w_{xx}\right)+\omega_{1}\left(\left|w\right|\right)+\omega_{2}\left(\right|w_{x}\left|\right)=0\text{ on }\mathcal{E},\\ w=0\text{ on }\partial\mathcal{E}.\end{cases}$
Note that 00 is a classical solution of the above equation, and thus also a (continuous) viscosity supersolution. Theorem 4.7 in Giga et al., (1991) then implies that $w\leq 0$ in $\mathcal{E}$, which yields the desired contradiction. This shows that $\mathcal{E}=\emptyset$ and, hence, that $v\geq u$ in $\bar{\mathcal{Y}}$.
The proof for the case in which $u$ is an $L^{p}$-supersolution, $v$ is a viscosity subsolution, and $u\geq v$ on $\partial\mathcal{Y}$ follows the same steps. $\blacksquare$
@@
Proposition 1 also holds when $F$ is degenerate elliptic (i.e, $\lambda=0$ in Assumption 4) since Theorem 4.7 in Giga et al., (1991) applies to degenerate elliptic operators. In this case, however, the existence of $L^{p}$-sub- or $L^{p}$-supersolutions is not guaranteed.
Proposition 1 and Lemma 1 yield the following corollary, which establishes the uniqueness of $L^{p}$-solutions.
@@
Suppose that Assumptions 4 and 8 hold and let $u$ and $v$ be, respectively, an $L^{p}$-subsolution [^13] The same observation regarding the absence of a constraint on the functions at the frontier $\partial\mathcal{Y}$ applies as in footnote 11. and an $L^{p}$-supersolution of
@@
If $v\geq u$ on $\partial\mathcal{Y}$, then $v\geq u$ on $\bar{\mathcal{Y}}$.
@@ 4.2 Stability Theorem
Stability theorems for both $L^{p}$-solutions and viscosity solutions of the Dirichlet problem (without obstacle) are well-known under a slight strengthening of the structure condition (SC). [^14] E.g., Theorem 6.1 in Crandall et al., (2000) gives a stability theorem when $F$ is Lipschitz continuous in $u_{x}$ and nonincreasing in $u$. We establish a more general stability theorem for our setting, which is then used to prove Theorem 1. This stability theorem is valid for a large class of nonlinear parabolic obstacle problems as long as $F$ satisfies the structure condition (SC).
@@
    for all $n\in\mathbb{N}$, $u^{n}=b^{n}$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}^{n}$, where the sequence $b^{n}$ is such that $b^{n}\to b$ pointwise for some function $b$ that is continuous on $\partial\mathcal{Y}$.
Let $F$ and $F^{n}$, $n\in\mathbb{N}$, be operators defined on $\mathcal{Y}\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}^{d}$ that satisfy Assumption 4 uniformly in $n$ and Assumption 5. Let $g^{n}$, $n\in\mathbb{N}$, be continuous functions on $\bar{\mathcal{Y}}$ such that $g^{n}\leq b^{n}$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}^{n}$.
Suppose that for any fixed $u_{xx}\in L^{p}\left(\mathcal{Y}\right)$, $F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. on $\mathcal{Y}$, and that either (i) $g^{n}\to g$ uniformly on compact subsets of $\mathcal{Y}$, or (ii) $g^{n}\uparrow g$ pointwise on $\mathcal{Y}$. If $u^{n}$, $n\in\mathbb{N}$, is an $L^{p}$-subsolution (supersolution) of
$\begin{cases}\max\left\{u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right),\,g^{n}-u^{n}\right\}=0\text{ on }\mathcal{Y}^{n},\\ u^{n}=b^{n}\text{ on }\partial\mathcal{Y}^{n},\end{cases}$ (Eqn)
@@
$\begin{cases}\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}=0\text{ on }\mathcal{Y},\\ u=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (Eq)
@@
By assumption, $u$ satisfies the regularity condition required to be an $L^{p}$-supersolution (requirement (i) in the definition). Moreover, since $u^{n}\geq b^{n}$ on $\partial\mathcal{Y}$ for all $n\in\mathbb{N}$, taking limits shows that $u\geq b$ on $\partial\mathcal{Y}$ and, hence, that $u$ satisfies the second requirement for being a supersolution.
We now prove the third condition, namely that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\leq 0\text{ a.e. on }\mathcal{Y}.$
Since $u^{n}$ is an $L^{p}$-supersolution, $g^{n}-u^{n}\leq 0$ on $\mathcal{Y}^{n}$. Moreover, $b^{n}\geq g^{n}$ on $\bar{\mathcal{Y}}\setminus\mathcal{Y}^{n}$ by assumption. So, for all $n\in\mathbb{N}$, $u^{n}\geq g^{n}$ on $\bar{\mathcal{Y}}$. Passing to the limit, $u\geq g$.
There remains to show that
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. on }{\mathcal{Y}},$
which will hold if
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. on }{\mathcal{O}}$
Thus, consider any strict bounded subset $\mathcal{O}$ of $\mathcal{Y}$. Since $\mathcal{Y}=\bigcup_{n\in\mathbb{N}}\mathcal{Y}^{n}$, compactness implies that there exists some integer $N\geq 1$ such that $\mathcal{Y}^{N}$ contains $\mathcal{O}$. Moreover, since the sequence $\{\mathcal{Y}^{n}\}_{n\geq 1}$ is increasing, $\mathcal{O}\subset\mathcal{Y}^{n}$ for all $n\geq N$. Without loss of generality, we assume that $N=1$. For all $m\in\mathbb{N}$, define
Since $u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)\leq 0$, we have for $n\geq m$ and almost every $(t,x)\in\mathcal{O}$,
$\underline{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\leq 0.$
Since $F^{n}$ satisfies Assumptions 4 and 5 for each $n$, we deduce that for all $m\in\mathbb{N}$ and all $v_{t}\in\mathbb{R}$, $\underline{H}^{m}(t,x,v_{t},\cdot)$ is Lipschitz continuous in $v_{xx}$ a.e. on $\mathcal{O}$ and
Moreover, for any $v_{xx}\in\mathbb{S}^{d}$, the structure condition (SC) implies that $\underline{H}^{m}$ is bounded in $L^{p}\left(\mathcal{O}\right)$ because the functions $u^{n}$, $n\in\mathbb{N}$, are uniformly bounded in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ as part of a weakly convergent sequence. Finally, by assumption, $u^{n}\rightharpoonup u$ in $W^{1,2,p}\left(\mathcal{O}\right)$. Theorem 4.2.6 in Krylov, (2018) then implies that, for all $m\in\mathbb{N}$,
$\underline{H}^{m}\left(t,x,u_{t},u_{xx}\right)=\underset{n\geq m}{\inf}\left\{u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\right\}=\underset{n\to\infty}{\lim}\,\underline{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\leq 0\text{ a.e. on }\mathcal{O}.$
Moreover, $u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. on $\mathcal{Y}$ by assumption. Letting $m\to\infty$, we obtain
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. on }\mathcal{O}.$
This shows that $u$ is an $L^{p}$-supersolution of (Eq).
Next, we prove that if, for each $n$, ${u^{n}}$ is an $L^{p}$-subsolution of (Eqn), then $u$ is an $L^{p}$-subsolution of (Eq).
We first note that in this case, we have $u^{n}\leq b^{n}$ on $\partial\mathcal{Y}$ for all $n\in\mathbb{N}$, . Taking limits, $u\leq b$ on $\partial\mathcal{Y}$.